# Are EU low-carbon structural funds efficient in reducing emissions?

### v2

@@ Abstract
We investigate the effectiveness of low-carbon expenditures from the European Structural and Investment Funds in reducing greenhouse gas emissions across EU regions. Using trend and cycle decomposition of two indicators – per capita emissions and emissions intensity – we find highly heterogeneous effects. In less developed regions, investments are linked to long-term increases in per capita emissions, while in transition and developed regions, effects are weak or not significant. Instrumental variable estimates confirm this pattern. When disaggregated by gas type, results remain inconsistent. Overall, our findings highlight the limited climate effectiveness of EU funding and the need for region-specific strategies.
@@ 1 Introduction
Consequently, we use the total amount spent on the low-carbon economy objective for the 2007-2013 and 2014-2020 budgeting periods as a measure of public climate investment. We then employ an econometric panel data model to investigate the influence of these investments on GHG emissions in European regions between 2007 and 2022 at the NUTS 2 level. Our analysis centres on two emission indicators: per capita emissions and emissions intensity (relative to regional gross domestic product (GDP)). We focus on the relationship between economic growth and environmental sustainability, and, to that end, we decompose the emission time series of each region into trend and cycle components. This allows us to determine expenditures’ short-term and long-term effects, considering that some regions have significantly advanced in terms of sustainability – achieving a given degree of decoupling – whereas many others have yet to do so. Our analysis also accounts for the direct impacts of policy by considering the OECD Environmental Policy Stringency (EPS) index (OECD, 2016) and average regional political opinions as covariates.
We find heterogeneous effects of low-carbon expenditures on GHG emissions across European regions. These impacts vary depending on the regional development level, investment period, and emission indicator. In the least developed regions, low-carbon expenditures are consistently associated with increases in long-term emissions per capita and, to a lesser extent, in emissions intensity. In contrast, results for developed and transition regions do not reveal significant reductions in long-term emissions trends, although there is some evidence of short-term mitigating effects, particularly for the 2007–2013 programme.
To address potential endogeneity, we implement a two-stage least squares (2SLS) model using lagged structural indicators and regional development categories as instruments. The instrumental variables estimates reinforce the main findings: low-carbon investments have heterogeneous effects across regions, with increases in per capita emissions in less developed regions and short-term reductions in developed regions. In transition regions, distinct patterns emerge that ultimately underscore the importance of the regional context.
When disaggregating emissions by gas type, we consistently find marked regional heterogeneity. The results reveal no significant reductions in emissions for developed regions, while in less developed regions, emissions – especially CO2 and N2O per capita – increase significantly. Effects in transition regions are mixed. Short-term dynamics also show variation by gas and region: investments under the 2007–2013 programme help reduce emissions volatility, particularly for CO2 in less developed and developed regions, and for N2O in transition regions.
@@ 2 Decoupling emissions: related literature
Yet a significant strand of literature doubts that the downward-sloping relationships between higher-income countries and pollution levels – which might imply a decoupling of economic activity with emissions – are enough to face climate change efficiently. Certainly, there are concerns about the possible shift in the production of pollution-intensive goods from developed to developing countries with less stringent environmental regulations (Hertwich and Peters, 2009; Levinson, 2023). This would weaken the Kuznets hypothesis since developed regions may reduce their at-home pollution levels by offshoring the production of environmentally harmful goods to nations where environmental regulations are less strict (Fischer and Heutel, 2013). While this points out that high-income countries should also look into reducing aggregate demand as a strategy for significant emissions reductions (Knight and Schor, 2014), it also requires the implementation of agreements at different scales: regional and global (Oberthür and Ott, 1999).
In this context, a growing body of literature indicates that international environmental agreements and climate policies are increasingly impacting emission trends, as reflected in the economic practices within countries, contributing to our understanding of the decoupling mechanism (Fischer and Heutel, 2013; Hildén et al., 2014; Haberl et al., 2020). Decoupling analyses aim to categorise the patterns of how emissions respond to economic growth – whether they increase at a slower rate, decrease, or stabilise as the economy grows (relative and absolute decoupling) or increase faster than the economy (coupling).
@@
Similarly, Doda (2014) finds that CO2 emissions are procyclical and more pronounced in countries with higher GDP per capita. Hence, the procyclicality also implies that economic growth without specific interventions often leads to increased emissions. Another fact is that emissions exhibit greater volatility compared to GDP, responding more dramatically to economic changes, and the extent to which emissions volatility exceeds GDP volatility decreases with the income level of the economies (Doda, 2014). These patterns are also observed at the subnational level: Dueñas and Mandel (2025), analysing NUTS 2 regions in Europe, further show that CO2 emissions growth rates follow an asymmetric and fat-tailed distribution and, similarly, that emissions volatility scales inversely with economic size of regions. This implies that more stable emission patterns through economic cycles in developed regions could be linked to more efficient energy use, a greater share of services in the economy, and potentially more effective environmental regulations and policies. In addition, emissions decrease more significantly during economic contractions than during expansions, suggesting an elastic response to economic downturns and inelasticity to economic growth (Sheldon, 2017).
@@
More directly related to the impacts of the EU Cohesion Policy, much of the literature has focused primarily on understanding its impact on regional development, specifically whether the funds have been effective in reducing regional disparities. For instance, very early on, Boldrin and Canova (2001) pointed out that these policies tend to be more redistributive than growth-oriented, reflecting political compromises over economic efficiency. However, investments in education and human capital, which have gained a significant portion of these funds, have shown medium-term positive impacts, emphasising the critical role of human capital development in fostering long-term regional growth and convergence (Rodríguez-Pose and Fratesi, 2004). Institutional quality plays a relevant role in the effectiveness of the funds. Evidence shown that local and regional governance directly affects economic growth and the effectiveness of structural funds (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014).
Although relatively few studies have explicitly focused on the effect of EU structural funds on reducing emissions, some anticipate that many of the same structural patterns that determine growth performance may also shape the trajectory of sustainable growth. Rodríguez-Pose and Bartalucci (2024) warn that while the green transition may benefit certain regions, it can also exacerbate existing territorial inequalities, particularly affecting less-developed regions. Nichifor et al. (2025) find that investments in renewables and environmental expenditures have positive effects on emissions reduction, but caution that these effects may exhibit diminishing returns beyond certain thresholds.
@@
In summary, in light of these findings, the effectiveness of the ESI funds in driving a low-carbon transition is conditional not only on structural factors such as institutional quality and regional development levels, but also on the phase of the economic cycle. Under Kuznets’ hypothesis, decoupling might appear problematic since regions would only begin to improve their environmental performance after reaching a certain level of wealth, implying a dependency between economic growth and environmental improvement. Therefore, seeking economic development without the corresponding environmental degradation is a great challenge for any policy framework.
Therefore, the central hypothesis of this paper is as follows:
@@
Hypothesis: Low‑carbon economy expenditures from the European Structural and Investment Funds have heterogeneous impacts on GHG emissions across regions at different stages of development, influencing both the long‑run trend and the magnitude of the business‑cycle component of emissions.
@@
To test this hypothesis, rather than focusing on total emissions, we concentrate on per capita emissions and emissions intensity (relative to regional GDP) . These indicators enhance comparability across regions by accounting for differences in population and economic scale, while also isolating environmental efficiency more effectively.
@@ 3.1 Data
For 2014-2020 raw categorisation data see: https://cohesiondata.ec.europa.eu/2014-2020-Categorisation/ESIF-2014-2020-ERDF-CF-ESF-raw-categorisation/xe4p-7b9q/about_data. These data contain the volume of EU support allocated to projects alongside a combination of categorisation dimension codes, including priority themes, the form of finance, the territorial context, location (using the Nomenclature of Territorial Units for Statistics, NUTS), and the economic dimension. We aggregate investments at NUTS level 2 (NUTS 2), as defined in version 2021. To do this, we consider all related investments, regardless of their original level of geographic specificity in the original database – be it national, NUTS-1, or even more granular at NUTS-3.
@@
For the 2014-2020 programme, the low-carbon expenditures are straightforward to determine since there is a categorical variable related to thematic objective 4 (TO-4), labelled as “low-carbon economy”. The aim of the expenditure can be tracked by the reported “intervention field” in the data. Specifically, related to TO-4, the intervention fields include renewable energy (wind/solar/biomass), energy efficiency renovation of public infrastructure, cycle tracks and footpaths, among others (see Table A1 in the Appendix).
For the 2007-2013 programme, there is no categorical variable explicitly associated with the low-carbon economy as in the 2014-2020 programme. However, the raw categorisation data describe the “priority” related to the reported expenditures, serving as a proxy for the aim of the expenditure. We use this information to create a low-carbon expenditures variable for this period. Priority labels related to this period include cycle tracks, intelligent transport systems, renewable energy (wind/solar/biomass/hydroelectric), energy efficiency, co-generation, energy management, and promotion of clean urban transport. The priority codes are reported in Table A1 of the Appendix.
A critical aspect of the ESI funds expenditure data is the absence of detailed annual expenditure information at the regional level. While reported expenditures can be easily tracked at the regional level, the reports correspond to the entire programming period. To our knowledge, there is no disaggregated time-varying consolidated official information available. Given this constraint, we built an annually growing cumulative investment series for each region and programme. This implies that, within each programming period, the total per capita expenditure is allocated progressively across the corresponding years. While this introduces a smoothing assumption, it allows us to preserve the annual variation in emissions and covariates. Therefore, considering the capital-intensive nature of the funded projects, we argue that their environmental effects are plausibly associated with the cumulative stock of investments. We further validate this modelling choice through robustness checks, including alternative cross-sectional estimations presented in Appendix Tables A2 and A3.
@@
Economic variables were obtained from the ARDECO (the Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy), which contains annual data at the regional level and is maintained and updated by the Joint Research Centre of the European Commission. We use variables such as the real gross domestic product, the population, and the real gross value added (GVA) for all industry sectors according to NACE Rev. 2 definitions. These sectors are A: Agriculture, Forestry and Fishing; B–E: Industry; F: Construction; G–J: Wholesale, Retail, Transport, Accommodation & Food Services, Information and Communication; K–N: Financial & Business Services; and O–U: Non-market Services. The choice of these variables is motivated by their relevance to understanding how economic activity is related to emissions. In particular, we include industry-level GVA to control direct sources of emissions and to better capture regional emission profiles, given that the participation of industries varies significantly across European regions.
@@ 3.2 Methodology
While our specification includes key relevant variables, we acknowledge that endogeneity may still pose concerns. For instance, one might argue that environmental regulations or improvements in energy efficiency could, in principle, influence regional output and investment decisions. For these reasons, we have included a five-year lag for all explanatory variables, ensuring that the predictors precede the observed emissions outcomes by a meaningful temporal distance. Therefore, while our specification mitigates (but does not fully eliminate) simultaneity concerns, it relies on the timing assumption that lagged expenditures impact emissions.
Moreover, the allocation of ESI funds is determined based on regional eligibility criteria related to development levels, making the investment variable exogenous to current emission levels. By interacting expenditures with regional development levels, we account for heterogeneous effects while controlling for potential confounding factors. Nonetheless, one could still argue that regions with higher emissions may have received more funds, representing a case of reverse causality or omitted variable bias. To test the robustness of our identification, we implement a 2SLS estimation using structural and institutional instruments plausibly uncorrelated with current emissions.
@@ 4.2 Effect in the emissions cycle
4.3 Instrumental variables estimations
To address potential endogeneity concerns related to low-carbon investments, we complement our baseline fixed-effects specification with an instrumental variables (2SLS) approach. While the use of lagged explanatory variables in the baseline models helps mitigate simultaneity bias, endogeneity may persist due to omitted variables or reverse causality. For instance, structurally more proactive regions might both attract higher low-carbon funding and implement complementary environmental measures not fully captured by our controls, potentially biasing the estimated coefficients.
We instrument per capita low-carbon expenditures from both programming periods (2007–2013 and 2014–2020) using regional structural indicators from the ARDECO database: (i) labour productivity per employee (log), (ii) consumption of fixed capital (log), (iii) hours worked per capita, and (iv) compensation of employees (log). These instruments are lagged by three years, not only to ensure temporal separation from emissions outcomes, but also based on the assumption that the allocation of funds is determined by pre-existing structural conditions. Additionally, we include regional development level dummies (TER and LDR, with DER as the reference category), which reflect the eligibility and allocation logic of the ESI funds. These variables are plausibly exogenous to short- and medium-term emissions dynamics.
First-stage results, presented in Table 4, confirm the strength of our instruments, with F-statistics exceeding conventional thresholds and statistically significant coefficients in the expected direction.
| Low Carbon Exp. pc07-13 | Low Carbon Exp. pc14-20 | |
| Labour productivity ppet-3 | -1.033*** | -0.683*** |
| (0.114) | (0.071) | |
| Hours worked per capitat-3 | 0.003*** | 0.001*** |
| (0.000) | (0.000) | |
| Compensation of Employeest-3 | -1.836*** | 0.422*** |
| (0.237) | (0.109) | |
| Consumption of fixed capitalt-3 | 1.990*** | -0.585*** |
| (0.237) | (0.110) | |
| TER | 0.970*** | 0.559*** |
| (0.117) | (0.066) | |
| LDR | 1.527*** | 1.329*** |
| (0.145) | (0.080) | |
| Time Dummies | Yes | Yes |
| F-statistic | 114.670 | 265.060 |
| Prob F | 0.000 | 0.000 |
| Observations | 1,694 | 1,687 |
| Notes: Standard errors in parentheses. Significance levels: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. TER and LDR refer to transition and less developed regions, respectively; DER is the reference group. | ||
For the second stage, we focus on the specifications that include the most comprehensive set of controls, those considering GVA disaggregated by sector. The column structure is intentionally aligned with the baseline results to facilitate direct comparison. Table 5 presents the second-stage results for both emissions performance indicators. Overall, the IV estimates reveal patterns that are qualitatively consistent with those obtained from the fixed-effects specification. To further support the validity of our instruments, we regress the second-stage residuals on the full set of instruments; the absence of significant coefficients suggests that the instruments are uncorrelated with the structural error term. Results of this test are reported in Appendix Table A4.
| Dependent variable: | ||||
| Trend | Cycle | |||
| (3t) | (6t) | (3c) | (6c) | |
| GHGEpc | GHGEi | GHGEpc | GHGEi | |
| Low Carbon Exp. pc | -0.587 | -0.224 | -0.601* | -0.144 |
| (0.393) | (0.389) | (0.345) | (0.201) | |
| Low Carbon Exp. pc | 5.692 | 4.207 | 0.460 | 0.340 |
| (3.468) | (3.847) | (0.815) | (0.846) | |
| Low Carbon Exp. pcTER | 1.399* | 1.949*** | -0.300 | 0.092 |
| (0.732) | (0.695) | (0.484) | (0.399) | |
| Low Carbon Exp. pcTER | -0.779 | 0.392 | -0.551 | -0.639 |
| (3.041) | (3.005) | (0.665) | (0.702) | |
| Low Carbon Exp. pcLDR | 1.126*** | 0.539 | 0.357 | 0.116 |
| (0.408) | (0.387) | (0.339) | (0.204) | |
| Low Carbon Exp. pcLDR | -2.350 | -2.487 | -0.556 | -0.360 |
| (3.189) | (3.462) | (0.740) | (0.772) | |
| EPSt-5 | 0.001 | -0.001 | 0.009*** | 0.011*** |
| (0.007) | (0.007) | (0.003) | (0.003) | |
| Political Ideologyt-5 | 0.006 | -0.010 | 0.003 | 0.002 |
| (0.010) | (0.013) | (0.003) | (0.003) | |
| GVAA,t-5 | 0.076** | 0.044 | -0.012 | -0.011 |
| (0.037) | (0.038) | (0.009) | (0.007) | |
| GVAB-E,t-5 | 0.213*** | -0.064 | -0.008 | -0.003 |
| (0.057) | (0.078) | (0.018) | (0.018) | |
| GVAF,t-5 | -0.110** | -0.134** | -0.046*** | 0.015 |
| (0.048) | (0.058) | (0.015) | (0.017) | |
| GVAG-J,t-5 | 0.205* | -0.129 | 0.018 | 0.078** |
| (0.113) | (0.129) | (0.035) | (0.032) | |
| GVAK-N,t-5 | -0.389*** | -0.436*** | -0.027 | 0.030 |
| (0.079) | (0.094) | (0.039) | (0.038) | |
| GVAO-U,t-5 | 0.147 | 0.303** | -0.180*** | -0.045 |
| (0.102) | (0.125) | (0.042) | (0.041) | |
| Observations | 3,408 | 3,408 | 3,408 | 3,408 |
| NUTS 2 fixed effects | Yes | Yes | Yes | Yes |
| Time fixed effects | Yes | Yes | Yes | Yes |
| Notes: Robust standard errors clustered at the NUTS 2 level are reported in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. Low Carbon Exp. pc07-13 and Low Carbon Exp. pc14-20 are instrumented. | ||||
For emissions per capita trends (Column 3t), low-carbon expenditures from the 2007–2013 programme are not statistically significant in developed regions, but the interaction term for less-developed regions remains positive and significant. This confirms that these regions experienced increases in long-term emissions per capita following higher investments. A key difference from the baseline is that the TER interaction becomes positive and significant, whereas it was previously negative and not significant. This suggests potential divergence in the response of transition regions under the IV approach. The 2014–2020 effects remain statistically not significant across development levels.
Regarding emissions intensity trends (Column 6t), we find a significant positive effect for the 2007-2013 expenditures in transition regions. This result departs from the baseline findings, where all interaction effects were not statistically significant.
Turning to the cyclical components, for emissions per capita (Column 3c), the 2007–2013 programme has a negative and statistically significant effect in developed regions, consistent with a smoothing effect on short-term volatility. However, the previously significant negative interaction effects for LDR are no longer significant under the IV specification. Similarly, for emissions intensity cycles (Column 6c), none of the coefficients related to expenditures are statistically significant, contrasting with the baseline results where some mitigating effects were observed.
In summary, the IV results reinforce the main conclusion of our baseline specification: low-carbon investments have different effects on emissions across different development levels. While some coefficients change in significance or sign, the overarching pattern, particularly the long-term positive association in less-developed regions, remains robust. These findings enhance confidence in the causal interpretation of the estimated effects and underscore the importance of considering regional structural differences when assessing climate policy outcomes.

### v3

@@ Abstract
We investigate the effectiveness of low-carbon expenditures from the European Structural and Investment Funds in reducing greenhouse gas emissions across EU regions. Using trend and cycle decomposition of per capita emissions and emissions intensity, along with a panel data approach that incorporates long lags to mitigate reverse causality, we find highly heterogeneous effects. In less developed regions, investments are associated with long-term increases in per capita emissions, whereas in transition and developed regions, the effects are weak or not significant. When disaggregated by gas type, results remain inconsistent. Our findings highlight that regional disparities challenge the effectiveness of EU climate efforts.
@@ 1 Introduction
Consequently, we use the total amount spent on the low-carbon economy objective for the 2007-2013 and 2014-2020 budgeting periods as a measure of public climate investment. We then employ an econometric panel data model to investigate the influence of these investments on GHG emissions in European regions between 2007 and 2022 at the NUTS 2 level. Our analysis centres on two emission indicators: per capita emissions and emissions intensity (relative to regional gross domestic product (GDP)). We focus on the relationship between economic growth and environmental sustainability, and, to that end, we decompose the emission time series of each region into trend and cycle components. This allows us to determine expenditures’ short-term and long-term effects, considering that some regions have significantly advanced in terms of sustainability—achieving a given degree of decoupling—whereas many others have yet to do so. Our analysis also accounts for the direct impacts of policy by considering the OECD Environmental Policy Stringency (EPS) index (OECD, 2016) and average regional political opinions as covariates.
To address potential endogeneity and reverse causality, we rely on a design with long lags of all right-hand-side variables, different specifications with controls including region-specific industry-composition (proxied by sectoral Gross Value Added (GVA)), and region and year fixed effects. We find heterogeneous effects of low-carbon expenditures on GHG emissions across European regions. These impacts vary depending on the regional development level, investment period, and emission indicator. In the least developed regions, low-carbon expenditures are consistently associated with increases in long-term emissions per capita and, to a lesser extent, in emissions intensity. In contrast, results for developed and transition regions do not reveal significant reductions in long-term emissions trends, although there is some evidence of short-term mitigating effects, particularly for the 2007–2013 programme.
When disaggregating emissions by gas type, we consistently find marked regional heterogeneity. The results reveal no significant reductions in emissions for developed regions, while in less developed regions, emissions—especially CO2 and N2O per capita—increase significantly. Effects in transition regions are mixed. Short-term dynamics also show variation by gas and region: investments under the 2007–2013 programme help reduce emissions volatility, particularly for CO2 in less developed and developed regions, and for N2O in transition regions.
@@ 2 Decoupling emissions: related literature
Yet a significant strand of literature doubts that the downward-sloping relationships between higher-income countries and pollution levels—which might imply a decoupling of economic activity with emissions—are enough to face climate change efficiently. Certainly, there are concerns about the possible shift in the production of pollution-intensive goods from developed to developing countries with less stringent environmental regulations (Hertwich and Peters, 2009; Levinson, 2023). This would weaken the Kuznets hypothesis since developed regions may reduce their at-home pollution levels by offshoring the production of environmentally harmful goods to nations where environmental regulations are less strict (Fischer and Heutel, 2013). While this points out that high-income countries should also look into reducing aggregate demand as a strategy for significant emissions reductions (Knight and Schor, 2014), it also requires the implementation of agreements at different scales: regional and global (Oberthür and Ott, 1999).
In this context, a growing body of literature indicates that international environmental agreements and climate policies are increasingly impacting emission trends, as reflected in the economic practices within countries, contributing to our understanding of the decoupling mechanism (Fischer and Heutel, 2013; Hildén et al., 2014; Haberl et al., 2020). Decoupling analyses aim to categorise the patterns of how emissions respond to economic growth—whether they increase at a slower rate, decrease, or stabilise as the economy grows (relative and absolute decoupling) or increase faster than the economy (coupling).
@@
Similarly, Doda (2014) finds that CO2 emissions are procyclical and more pronounced in countries with higher GDP per capita. Hence, the procyclicality also implies that economic growth without specific interventions often leads to increased emissions. Another fact is that emissions exhibit greater volatility compared to GDP, responding more dramatically to economic changes, and the extent to which emissions volatility exceeds GDP volatility decreases with the income level of the economies (Doda, 2014). These patterns are also observed at the subnational level: Dueñas and Mandel (2025), analysing NUTS 2 regions in Europe, further show that CO2 emissions growth rates follow an asymmetric and fat-tailed distribution and, similarly, that emissions volatility scales inversely with the economic size of regions. This implies that more stable emission patterns through economic cycles in developed regions could be linked to more efficient energy use, a greater share of services in the economy, and potentially more effective environmental regulations and policies. In addition, emissions decrease more significantly during economic contractions than during expansions, suggesting an elastic response to economic downturns and inelasticity to economic growth (Sheldon, 2017).
@@
While the above studies highlight the cyclical and long-term dynamics of emissions in response to policy, the Cohesion Policy literature has concentrated mainly on regional development objectives, which is its primary objective. Much of the literature has focused primarily on understanding its impact on regional development, specifically whether the funds have been effective in reducing regional disparities. For instance, very early on, Boldrin and Canova (2001) pointed out that these policies tend to be more redistributive than growth-oriented, reflecting political compromises over economic efficiency. However, investments in education and human capital, which have gained a significant portion of these funds, have shown medium-term positive impacts, emphasising the critical role of human capital development in fostering long-term regional growth and convergence (Rodríguez-Pose and Fratesi, 2004). Institutional quality plays a relevant role in the effectiveness of the funds. Evidence shows that local and regional governance directly affects economic growth and the effectiveness of structural funds (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014).
Although relatively few studies have explicitly focused on the effect of EU structural funds on reducing emissions, some anticipate that many of the same structural patterns that determine growth performance may also shape the trajectory of sustainable growth. In particular, some contributions emphasise the importance of regional asymmetries in the green transition. Rodríguez-Pose and Bartalucci (2024) warn that while the green transition may benefit certain regions, it can also exacerbate existing territorial inequalities, particularly affecting less-developed regions. They claim that less-developed, peri-urban, and rural regions in Southern and Eastern Europe are systematically more vulnerable, while metropolitan areas are better positioned to embrace green opportunities thanks to stronger infrastructure, skills, and institutional capacity. At the same time, the micro-foundations of the green upgrading further stress the role of regional capability-relatedness (Boschma et al., 2017). Santoalha and Boschma (2021) show that regions are more likely to diversify into green technologies when related capabilities are present. They also claim that political support does not exert a strong direct effect but interacts with capabilities: regional-level support strengthens, while national-level support may moderate the role of relatedness. Consistent with this territorial heterogeneity, Dueñas and Mandel (2025) show that European regions experience markedly different emission dynamics under the EU ETS, with less-developed regions displaying higher volatility and more uncertain decarbonisation pathways than their richer counterparts, suggesting that the growth–emissions nexus is mediated by regional characteristics, institutions, and historical trajectories. Complementing this evidence, Nichifor et al. (2025) find that in Eastern Europe, renewable investments and environmental expenditures reduce emissions on average, but their effectiveness is spatially uneven and may exhibit diminishing returns beyond policy-relevant thresholds, underscoring the need for coordinated, place-based strategies.
@@
In summary, in light of these findings, the effectiveness of the ESI funds in driving a low-carbon transition is conditional not only on structural factors such as institutional quality and regional development levels, but also on the phase of the economic cycle. Less-developed regions, characterised by higher vulnerability and weaker absorptive capacity, may struggle to convert expenditures into immediate emission reductions, potentially exhibiting weaker or delayed effects (and even short-run increases), while more developed regions are better positioned to leverage such funds into abatement. These asymmetries align with the logic of the environmental Kuznets hypothesis, suggesting that decoupling may be problematic, as regions tend to improve their environmental performance only after reaching a certain level of wealth. Therefore, seeking economic development while reducing disparities across European regions, and at the same time avoiding environmental degradation, remains a major challenge in the EU policy framework.
Accordingly, the central hypothesis of this paper is as follows:
@@
Hypothesis 1: Low‑carbon economy expenditures from the European Structural and Investment Funds have heterogeneous impacts on GHG emissions across regions at different stages of development, influencing both the long‑run trend and the magnitude of the business‑cycle component of emissions.
Building on recent contributions emphasising regional asymmetries of European regions (as in Santoalha and Boschma 2021 and Rodríguez-Pose and Bartalucci 2024), it can be further expected the following:
Hypothesis 2: The effectiveness of low-carbon expenditures is conditional on regional development levels, with weaker—and in some cases emission-increasing—impacts in less-developed regions.
@@
To test these hypotheses, rather than focusing on total emissions, we concentrate on per capita emissions and emissions intensity (relative to regional GDP) . These indicators enhance comparability across regions by accounting for differences in population and economic scale, while also isolating environmental efficiency more effectively.
@@ 3.1 Data
For 2014-2020 raw categorisation data see: https://cohesiondata.ec.europa.eu/2014-2020-Categorisation/ESIF-2014-2020-ERDF-CF-ESF-raw-categorisation/xe4p-7b9q/about_data. These data contain the volume of EU support allocated to projects alongside a combination of categorisation dimension codes, including priority themes, the form of finance, the territorial context, location (using the Nomenclature of Territorial Units for Statistics, NUTS), and the economic dimension. We aggregate investments at NUTS level 2 (NUTS 2), as defined in version 2021. To do this, we consider all related investments, regardless of their original level of geographic specificity in the original database—be it national, NUTS-1, or even more granular at NUTS-3.
@@
For the 2014-2020 programme, the low-carbon expenditures are straightforward to determine since there is a categorical variable related to thematic objective 4 (TO-4), labelled as “low-carbon economy”. The aim of the expenditure can be tracked by the reported “intervention field” in the data. Specifically, related to TO-4, the intervention fields include renewable energy (wind/solar/biomass), energy efficiency renovation of public infrastructure, cycle tracks and footpaths, among others (see Appendix A1).
For the 2007-2013 programme, there is no categorical variable explicitly associated with the low-carbon economy as in the 2014-2020 programme. However, the raw categorisation data describe the “priority” related to the reported expenditures, serving as a proxy for the aim of the expenditure. We use this information to create a low-carbon expenditures variable for this period. Priority labels related to this period include cycle tracks, intelligent transport systems, renewable energy (wind/solar/biomass/hydroelectric), energy efficiency, co-generation, energy management, and promotion of clean urban transport. The priority codes are reported in Appendix A1.
A critical aspect of the ESI funds expenditure data is the absence of detailed annual expenditure information at the regional level. While reported expenditures can be easily tracked at the regional level, the reports correspond to the entire programming period. To our knowledge, there is no disaggregated time-varying consolidated official information available. Given this constraint, we built an annually growing cumulative investment series for each region and programme. This implies that, within each programming period, the total per capita expenditure is allocated progressively across the corresponding years. While this introduces a smoothing assumption, it allows us to preserve the annual variation in emissions and covariates. Therefore, considering the capital-intensive nature of the funded projects, we argue that their environmental effects are plausibly associated with the cumulative stock of investments. We further validate this modelling choice through robustness checks, including alternative cross-sectional estimations presented in Appendix A2.
@@
Economic variables were obtained from the ARDECO (the Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy), which contains annual data at the regional level and is maintained and updated by the Joint Research Centre of the European Commission. We use variables such as the real gross domestic product, the population, and the real gross value added for all industry sectors according to NACE Rev. 2 definitions. These sectors are A: Agriculture, Forestry and Fishing; B–E: Industry; F: Construction; G–J: Wholesale, Retail, Transport, Accommodation & Food Services, Information and Communication; K–N: Financial & Business Services; and O–U: Non-market Services. The choice of these variables is motivated by their relevance to understanding how economic activity is related to emissions. In particular, we include industry-level GVA to control direct sources of emissions and to better capture regional emission profiles, given that the participation of industries varies significantly across European regions.
@@ 3.2 Methodology
While our specification includes key relevant variables, we acknowledge that endogeneity may still pose concerns. For instance, one might argue that environmental regulations or improvements in energy efficiency could, in principle, influence regional output and investment decisions. Moreover, the allocation of ESI funds is determined based on regional eligibility criteria related to development levels, making the investment variable exogenous to current emission levels. By interacting expenditures with regional development levels, we account for heterogeneous effects while controlling for potential confounding factors. Nonetheless, one could still argue that regions with higher emissions may have received more funds, representing a case of reverse causality or omitted variable bias.
For these reasons, we have included a five-year lag for all explanatory variables, ensuring that the predictors precede the observed emissions outcomes by a meaningful temporal distance. Therefore, our specification mitigates—though does not fully eliminate—simultaneity concerns and relies on the timing assumption that lagged expenditures impact emissions. A plausible strategy would have been to implement instrumental variables; however, given the nature of the impact of low-carbon investments—which take time to materialise—it is difficult to identify strong instruments.888As a robustness check, in Appendix A3, we implemented 2SLS estimations using structural and institutional variables (e.g. labour productivity, hours worked per capita, consumption of fixed capital, compensation of employees) as potential instruments. Despite these instruments proved to be valid, they are not strong enough to support a reliable specification, specifically the Kleibergen-Paap rk Wald F statistic is well below conventional thresholds. Thus, these IV results must be interpreted with caution.
@@ 4.1 Effect on emissions trend
These results support H1, as the effects of low-carbon expenditures are heterogeneous across development levels and emission indicators. Moreover, the consistently positive and significant interactions for LDR alongside weak or null effects for DER/TER provide evidence in favour of H2.
@@ 4.2 Effect in the emissions cycle
These findings reinforce H1 and H2, showing that low-carbon investments affect not only long-run trajectories but also the business-cycle component of emissions, with short-run impacts that vary heterogeneously across regions depending on their level of development.